# Why Larp?! A Synthesis Paper on Live Action Roleplay in Relation to HCI Research and Practice

### v1

@@ 6.1. Larp: The Object
The Object was designed within a research collaboration project together with artists, designers, museum professionals and researchers to support the creation of museum experiences. It was included in a collection of similar scenarios, all created to provide a sensitising and team-building experience for design teams working in a museum context, to reflect upon and learn about relevant concepts and theories related to their practice. The Object focuses on how we value objects, the life cycle of museum objects, and on what stories museums choose to tell. The script includes pre-workshop instructions, the larp itself, and suggestions for a structured debrief. The script is short but detailed, written to be accessible to design teams without access to expert larp facilitators. The Object was developed by three members of a research group, all with previous experience as larpers and larp designers.
6.1.1. Basic facts about the larp
Larp Designers: Karin Johansson, Jon Back, Annika Waern
Type of venue: an ordinary room
Number of participants: 4-9
Duration: 1 hour and 20 min (including workshop and debrief)
Props/costumes: An object such as a small flower pot. Printed out roles and instructions.
Target group: Museum personnel/ design teams working with museums
Purpose/learning goals: reflecting and discussing provenance, value, stories and life-cycles of museum objects, what stories we tell at museums and what alternate stories there could be. New museology, critical perspectives, experience design and use of technology in cultural heritage practices.
Short summary of plot/story: Through small scenes the participants follow an object from its creation, it being collected for a museum, it being exhibited and finally discarded from the museum collection. The participants play different roles in different scenes, for example villagers or museum personnel, and the roles have different agendas and attitudes.
@@ 6.2. Larp: The Deadline
The Deadline is one of five scenarios developed for an interdisciplinary research curation center within a university. The center helps interdisciplinary research teams develop grant proposals and learn how to collaborate. The purpose of the scenario is to guide researchers through situations of escalated conflict that often arise during stressful situations. In the scenario, the characters are researchers in conflict related to whose research contributions are most relevant and valuable. The goal of the larp is to practice nonviolent communication skills such as “I-statements” within a professional environment, allow researchers to think beyond their positions and identify not just their interests, but also feelings and basic human needs underlying the conflict.
6.2.1. Basic facts about the larp
Larp Designers: Sarah Lynne Bowman, Josefin Westborg, and Kaya Toft Thejls
Type of venue: an ordinary meeting room
Number of participants: 3-5 + facilitator character
Duration: 3 hours
Props/costumes: None
Target group: research scholars seeking funding in interdisciplinary collaborations
Purpose/learning goals: 1: Practice receiving and responding to feedback that might be upsetting or frustrating. 2: Explain one’s research agenda in the hopes that others from different disciplines can understand and value it. 3: Persuade others to address one’s interests within a research group through self-advocacy. 4: Articulate interests, needs, and feelings through the use of nonviolent communication. 5: Collaborate to search for win-win scenarios in order to complete the proposal successfully.
Short summary of plot/story: The scenario is set during the development process of a research grant proposal. A group of multidisciplinary academics are trying to conduct revisions on their proposal based on each of their interests and research agendas. Conflicts emerge in the group that must be addressed or the proposal will not meet the deadline. The purpose of the scenario is to help researchers address the needs of everyone in a group most peacefully and effectively through the practices of conflict transformation (Lederach, 2014; Gullick et al., 2020; Rosenberg and Chopra, 2015).
@@ 6.3. Larp: The Anywear Academy
This larp was designed as part of a nationally funded project aimed at broadening interest and participation in STEM topics and careers. The purpose of this larp is to let middle-school girls familiarise themselves with technology design and development in ways that encourage building computational and design skills, interest, and self-efficacy. A paper published in DIS 2022 covers the concept, design, and impact of the first iteration of the larp (Fey et al., 2022), while a series of publications are currently either under review or in progress. Additionally, the larp material is publically available for use  [^36] https://www.anywear-academy.org/.
Figure 8. Photos from the social wearables edu-larp (SWEL) camp for middle school girls.
6.3.1. Basic facts about the larp
Larp Designers: James Fey, Katherine Isbister, Ella Dagan, Raquel Robinson, The Game Academy
Type of venue: multiple rooms, one for crafting/coding, and another for the missions
Number of participants: 7-15
Duration: 5-day camp
Props/costumes: Fairy-inspired (crowns, vines, tutus, fairy wings), superhero-inspired (capes, masks), all costumes/wearables designed by campers.
Purpose/learning goals: a larp-based summer camp in which middle school age girls create social wearables, toward building computational and design skills, as well as interest in and self-efficacy about coding and design.
Short summary of plot/story: Campers participate in three different mission chains, or plots, separated by theme and setting. Campers are trainees of the ‘Anywear Academy’, tasked with traveling undercover to different worlds. There are three worlds the campers travel to: Elphame, a fairy-themed magical world; Metro City, a superhero-themed world, and Earhart station, a space station. Before every mission, an NPC (non-player character–e.g. a camp staff member playing a role) conducts a short briefing, explaining the mission objectives. Working from this briefing, campers design wearables that allowed them to blend into each setting, as well as meet specific technical challenges laid out to them. In Elphame, campers have magic powers using their wearables, but they are tasked to disguise them so they appear like magic rather than a tech device (which fairies mistrust). In Metro City, campers create their own superhero personas with their wearables giving them special superpowers, to investigate a new villain that has started to plague the city. In Earhart station, campers must bring the space station back online using their wearables for communication between one another, as they are not able to speak due to the air-support systems being offline.
@@ 6.4. Larp: True Colors Embedded in New Gyr
This example is a Research-through-Design investigation of how to design wearables to support pro-social behavior. Academic researchers worked closely with a local larp community to design and deploy wearable technology in one instance of a larp series exploring life in a fictional galaxy. The wearable was been demoed as part of UbiComp/ISWC’19 (Dagan et al., 2019a). Key design research insights from this larp were presented as part of a CHI’19 paper (Dagan et al., 2019c). The work was also paved the way for intermediate-level knowledge (Höök and Löwgren, 2012), giving rise to a design framework for social wearables, and published at DIS’19 (Dagan et al., 2019d), and in an MIT Book (Buruk et al., 2023).
6.4.1. Basic facts about the larp
Larp Designers: Ella Dagan, Katherine Isbister, Elena Márquez Segura, James Fey
Type of venue: A whole conference and retreat center, with 3 buildings, multiple rooms, and outdoor spaces
Number of participants: 109 (91 players; NPCs/staff: 15; 3 design researchers)
Duration: 3 days (approx. 32 hours of gameplay)
Props/costumes: Sci-Fi. Designed, and built by participants. Augment characters wore a wearable designed for the purpose of the design research project. Target group: General population ¿18 years old
Purpose/learning goals: the larp was meant to be a sandbox for players to explore and experiment, with minimal organizer-introduced plot elements, rich in player-driven and player-organized plot(s).
Short summary of plot/story: On a fictional planet several ”human variants” coexist, not always in peace: Humans; Augments (Humans augmented with technology); Evos (humans with altered genome); and Androids (artificial intelligent agents). The Augments (13 players) were humans who had resorted to technological augmentation to improve their bodies, for practical reasons “such as enhancing physical strength, expanding memory, or removing the need to breathe.” Their cultural status was low-class, taking on physical labor or working in hazardous environments. The research group worked together with the larp designers to develop digital wearables for the Augments, designed to foster pro-social behavour within the group and with characters from the other factions.
@@ 7.0.1. Research into larps
Larp is a cultural expression in its own right. It is also a global phenomenon with many million players all over the world, mainly grassroot organized via social media and other digital platforms. Tech has increasingly been integrated in different ways into these larps, as has been presented in the practitioner overview. Despite all this, little has been published, as is pointed out in the literature review, within HCI research about existing larps practices. And when larps are studied, the research focus has largely been on experiments with innovative technology, rather than established practices of technology use in larp, that the practitioner overview shows is abundant. This also goes for larp as relevant interaction design practice. The consequence is that we know too little about how larps are designed, and what considerations are important for the use of technology within them. Understanding larp design better could inform other types of experience and/or activity design, especially in relation to co-creation, first person perspectives and in the creation of experiences that are at the same time individualized and collective, given the social and collective nature of larp experiences.
@@ 7.0.2. Research for larps
Hand-crafted novel technologies and existing digital tools are already used in larps, as is shown in the practitioner overview, and larp designers and players use practical knowledge to build larps, costumes, and settings. This practice remains under-researched, and could potentially inform design in adjacent domains. Furthermore, many larp designs require facilitators to be physically present and active at all times, which is why many larps become confined to specific spaces or small numbers of players. Technology solutions could support such larps in scaling up or becoming more easily re-stageable. In the literature review, this type of research seems scarce, even though, as shown previously in this paper, larp is a vibrant and large subculture so there could be interesting to support and inform larp design processes, and to for instance find technology solutions that could be key in supporting a wider use of larp for serious purposes such as training or therapy. Finally, even though technology-related HCI research has been dominated by technology innovation, there is still room for innovation in the use of technology for creating new larp experiences.
@@ 7.0.3. Research through larps
Several publications have already utilised larp as a tool to explore relevant HCI topics, as is shown in the literature review and through the larp research exemplars. One reason for this is the versatile format that larp offers. In particular, larp has a role as a design tool that can be used throughout the whole design process (Márquez Segura et al., 2019) to sensitize designers and other stakeholders to particularly sensitive topics, situations, and scenarios; to come up with interesting technology design ideas (e.g. (Fey et al., 2022)); and to test, evaluate and even iterate technology designs and prototypes (e.g. (Márquez Segura et al., 2018; Dagan et al., 2019c)). Larps also align well with critical approaches such as design fiction; they are particularly well suited to imagine futures and future designs and critically experience and reflect on them and their impact. Futuring was identified as a reoccurring topic in both the literature review, as well as seen occurring though the practitioner overview. Redström, in his book Making Design Theory, argues that design research could and should engage with futuring: ”Design research addressing matters such as attitudes, values, and behaviours, does not necessarily depend on future technologies or opportunities to be possible. On the contrary, it depends on questioning what is now, asking what if things were already different” (Redström, 2017). Larp, as mentioned in this paper, is about making a mental journey, putting oneself in an immersive situation and through transformative play imagining different realities. This gives unique opportunities to engage in futuring, independent of future technologies, since we can pretend through the larp that they already exist, and then explore what that would mean for a group of people, part of the society, or for our own behaviors and emotions. How this can be done can be seen for instance by the True Color exemplar. Further, larps can be suitable when working with value-based design research, such as value sensitive design, socially responsible design, reflective design, and design for social sustainability, as it is a good way of shifting perspectives, sparking discussions and getting emotional understanding for different concepts and situations.
@@ 7.1. Benefits of Utilising Larp in HCI
  • Playification - larp offers many of the benefits of gamification but without strong extrinsic motivations such as badges, scores and leaderboards. Instead, larp motivates through playful exploration and co-creative storytelling.
@@ 7.2. Challenges with Utilising Larp in HCI Research
There are challenges with utilising larp in HCI research, that need to be thought through and handled, as well as reflected upon when making informed choices, both when deciding if larp is to be utilised, as well as how. Some of those challenges are more general when designing and facilitating larps, and there is already an abundance of literature on this (Koljonen et al., 2019; Back, 2014; Duus Henriksen et al., 2011), but there are also specific challenges for utilising larp in HCI relevant research, and this is the focus of this section.
@@ 7.3. Handling the Challenges
There is a need to develop good methods and research practices that are adapted to the design logic of larps. In this, it is vital to learn from fields such as education research, ethnography and anthropology, on how to stage and study situated experiences.
@@
We propose discussions within the HCI community, to find new methods for documentation, and develop formats for rigorous mediation of research in relation to larp. There are several practical examples available in the literature that can inform how to structure methods and formats of larps (Waern et al., 2020; Fey et al., 2022), but future research focusing specifically on developing such methods could be helpful for the research community. This discussion needs to also include how design research focused on larp can and should be presented, as the paper format can be ill-suited for situated and embodied experiences such as larp.
@@ 8. Conclusion
  • Larp is for several reasons relevant for the CHI community

### v2

@@ 6.1. Larp: The Object
The Object was designed within a research collaboration project together with artists, designers, museum professionals and researchers to support the creation of museum experiences. It was included in a collection of similar scenarios, all created to provide a sensitising and team-building experience for design teams working in a museum context, to reflect upon and learn about relevant concepts and theories related to their practice. The Object focuses on how we value objects, the life cycle of museum objects, and on what stories museums choose to tell. The script includes pre-workshop instructions, the larp itself, and suggestions for a structured debrief. The script is short but detailed, written to be accessible to design teams without access to expert larp facilitators. The Object was developed by three members of a research group, all with previous experience as larpers and larp designers. Basic facts about the larp can be found in Table 1.
Table 1. Basic facts about the larp: The Object
Name of larp The Object
Number of players 4-9 players, including facilitator
Venue Ordinary room
Duration 1 hour 20 min
Target group Museum design teams
Props An object such as a small flower pot. Printed out roles and instructions.
Short description
Through small scenes the participants follow an object from its creation,
it being collected for a museum, it being exhibited and finally discarded
from the museum collection. The participants play different roles in
different scenes, for example villagers or museum personnel, and the
roles have different agendas and attitudes.
Purpose of the larp
Sensitize design teams at museums towards concepts such as cultural
heritage, cultural appropriation and deaccession. Reflecting and
discussing provenance, value, stories and life-cycles of museum
objects, what stories we tell at museums and what alternate stories
there could be. New museology, critical perspectives, experience design
and use of technology in cultural heritage practices.
Larp Designers Karin Johansson, Jon Back, Annika Waern
@@ 6.2. Larp: The Deadline
The Deadline is one of five scenarios developed for an interdisciplinary research curation center within a university. The center helps interdisciplinary research teams develop grant proposals and learn how to collaborate. The purpose of the scenario is to guide researchers through situations of escalated conflict that often arise during stressful situations. In the scenario, the characters are researchers in conflict related to whose research contributions are most relevant and valuable. The goal of the larp is to practice nonviolent communication skills such as “I-statements” within a professional environment, allow researchers to think beyond their positions and identify not just their interests, but also feelings and basic human needs underlying the conflict. Basic facts about the larp can be found in Table 2.
Table 2. Basic facts about the larp: The Deadline
Name of larp The Deadline
Number of players 3-5 + facilitator character
Venue Ordinary room
Duration 3 hours
Target group Research scholars seeking funding in interdisciplinary collaborations
Props None
Short description
The scenario is set during the development process of a research grant
proposal. A group of multidisciplinary academics are trying to conduct
revisions on their proposal based on each of their interests and
research agendas. Conflicts emerge in the group that must be addressed
or the proposal will not meet the deadline. The purpose of the scenario
is to help researchers address the needs of everyone in a group most
peacefully and effectively through the practices of conflict
transformation (Lederach, 2014; Gullick et al., 2020; Rosenberg and Chopra, 2015)
Purpose of the larp
1) Practice receiving and responding to feedback that might be
upsetting or frustrating.
2) Explain one’s research agenda in the hopes that others from different
disciplines can understand and value it.
3) Persuade others to address one’s interests within a research group
through self-advocacy.
4) Articulate interests, needs, and feelings through the use of
nonviolent communication.
5) Collaborate to search for win-win scenarios in order to complete
the proposal
successfully.
Larp Designers Sarah Lynne Bowman, Josefin Westborg, and Kaya Toft Thejls
(Lederach, 2014; Gullick et al., 2020; Rosenberg and Chopra, 2015).
@@ 6.3. Larp: The Anywear Academy
This larp was designed as part of a nationally funded project aimed at broadening interest and participation in STEM topics and careers. The purpose of this larp is to let middle-school girls familiarise themselves with technology design and development in ways that encourage building computational and design skills, interest, and self-efficacy. A paper published in DIS 2022 covers the concept, design, and impact of the first iteration of the larp (Fey et al., 2022), while a series of publications are currently either under review or in progress. Additionally, the larp material is publically available for use  [^21] https://www.anywear-academy.org/. Basic facts about the larp can be found in Table 3.
Figure 10. Photos from The Anywear Academy 5-day summer camp for middle school girls.
Table 3. Basic facts about the larp: The Anywear Academy
Name of larp The Anywear Academy
Number of players 7-15
Venue Multiple rooms, one for crafting/coding, and another for the missions
Duration 5-day camp
Target group Research scholars seeking funding in interdisciplinary collaborations
Props
1) Fairy-inspired items (crowns, vines tutus, fairy wings)
2) Superhero-inspired items (capes, masks).
All costumes/wearables designed by campers.
Short description
Campers participate in three different mission chains, or plots,
separated by theme and setting. Campers are trainees of the
‘Anywear Academy’, tasked with traveling undercover to different
worlds. There are three worlds the campers travel to: Elphame, a
fairy-themed magical world; Metro City, a superhero-themed world,
and Earhart station, a space station. Before every mission, an NPC
(non-player character–e.g. a camp staff member playing a role)
conducts a short briefing, explaining the mission objectives. Working
from this briefing, campers design wearables that allowed them to
blend into each setting, as well as meet specific technical challenges
laid out to them.
Purpose of the larp
A larp-based summer camp in which middle school-age girls create
social wearables, toward building computational and design skills,
as well as interest in and self-efficacy about coding and design.
Larp Designers
James Fey, Katherine Isbister, Ella Dagan, The Game Academy,
Raquel Robinson, Selin Ovali
@@ 6.4. Larp: True Colors Embedded in New Gyr
This example is a Research-through-Design investigation of how to design wearables to support pro-social behavior. Academic researchers worked closely with a local larp community to design and deploy wearable technology in one instance of a larp series called New Gyr, exploring life in a fictional galaxy. The wearable was been demoed as part of UbiComp/ISWC’19 (Dagan et al., 2019a). Key design research insights from this larp were presented as part of a CHI’19 paper (Dagan et al., 2019c). The work was also paved the way for intermediate-level knowledge (Höök and Löwgren, 2012), giving rise to a design framework for social wearables, and published at DIS’19 (Dagan et al., 2019d), and in an MIT Book (Buruk et al., 2023). Basic facts about the larp can be found in Table 4.
Table 4. Basic facts about the larp: New Gyr
Name of larp New Gyr (’True Colors’ tech was embedded within the larp)
Number of players 109 (91 players; NPCs/staff: 15; 3 design researchers)
Venue
Conference/retreat center with 3 buildings, multiple rooms and
outdoor spaces
Duration 3 days (approx. 32 hours of gameplay)
Target group General population ¿18 years old
Props
Sci-Fi themed props and costumes, all designed and built by
participants. Augment characters wore a wearable designed for
the purpose of the design research project.
Short description
On a fictional planet several ”human variants” coexist, not always
in peace: Humans; Augments (Humans augmented with technology);
Evos (humans with altered genome); and Androids (artificial
intelligent agents). The Augments (13 players) were humans who
had resorted to technological augmentation to improve their
bodies, for practical reasons “such as enhancing physical strength,
expanding memory, or removing the need to breathe.” Their
cultural status was low-class, taking on physical labor or working
in hazardous environments. The research group worked together
with the larp designers to develop digital wearables for the
Augments, designed to foster pro-social behaviour within the group
and with characters from the other factions.
Purpose of the larp
The larp was meant to be a sandbox for players to explore and
experiment, with minimal organizer-introduced plot elements,
rich in player-driven and player-organized plot(s).
Larp (Tech) Designers Ella Dagan, Katherine Isbister, Elena Marquez Segura, James Fey
@@ 7.0.1. Research into larps
Larp is a cultural expression in its own right. It is also a global phenomenon with many million players all over the world, mainly grassroot organized via social media and other digital platforms. Tech has increasingly been integrated in different ways into these larps, as has been presented in the practitioner overview. Despite all this, little has been published, as is pointed out in the literature review, within HCI research about existing larps practices. And when larps are studied, the research focus has largely been on experiments with innovative technology, rather than established practices of technology use in larp, that the practitioner overview shows is abundant. This also goes for larp as relevant interaction design practice. The consequence is that we know too little about how larps are designed, and what considerations are important for the use of technology within them. Understanding larp design better could inform other types of experience and/or activity design, especially in relation to co-creation, first person perspectives and in the creation of experiences that are at the same time individualized and collective, given the social and collective nature of larp experiences. An example of a research project engaging with research into larp could be a comparative study between an online larp and an on-site larp, to identify what interactions are shared between the two types of larp, and to what extent for instance the embodiment and tangibility of a physical version of a larp affects the experience.
@@ 7.0.2. Research for larps
Hand-crafted novel technologies and existing digital tools are already used in larps, as is shown in the practitioner overview, and larp designers and players use practical knowledge to build larps, costumes, and settings. This practice remains under-researched, and could potentially inform design in adjacent domains. Furthermore, many larp designs require facilitators to be physically present and active at all times, which is why many larps become confined to specific spaces or small numbers of players. Technology solutions could support such larps in scaling up or becoming more easily re-stageable. In the literature review, this type of research seems scarce, even though, as shown previously in this paper, larp is a vibrant and large subculture so there could be interesting to support and inform larp design processes, and to for instance find technology solutions that could be key in supporting a wider use of larp for serious purposes such as training or therapy. Finally, even though technology-related HCI research has been dominated by technology innovation, there is still room for innovation in the use of technology for creating new larp experiences. An example of a research project engaging with research for larp could be a Research through Design project developing a diegetic magical looking glass through which different larp groups could gain secret knowledge about other groups, and then try and tweak the design to create maximum leverage from a larp interaction perspective, identifying which features of such a design that would best promote immersion and a better larp experience.
@@ 7.0.3. Research through larps
Several publications have already utilised larp as a tool to explore relevant HCI topics, as is shown in the literature review and through the larp research exemplars. One reason for this is the versatile format that larp offers. In particular, larp has a role as a design tool that can be used throughout the whole design process (Márquez Segura et al., 2019) to sensitize designers and other stakeholders to particularly sensitive topics, situations, and scenarios; to come up with interesting technology design ideas (e.g. (Fey et al., 2022)); and to test, evaluate and even iterate technology designs and prototypes (e.g. (Márquez Segura et al., 2018; Dagan et al., 2019c)). Larps also align well with critical approaches such as design fiction; they are particularly well suited to imagine futures and future designs and critically experience and reflect on them and their impact. Futuring was identified as a reoccurring topic in both the literature review, as well as seen occurring though the practitioner overview. Redström, in his book Making Design Theory, argues that design research could and should engage with futuring: ”Design research addressing matters such as attitudes, values, and behaviours, does not necessarily depend on future technologies or opportunities to be possible. On the contrary, it depends on questioning what is now, asking what if things were already different” (Redström, 2017). Larp, as mentioned in this paper, is about making a mental journey, putting oneself in an immersive situation and through transformative play imagining different realities. This gives unique opportunities to engage in futuring, independent of future technologies, since we can pretend through the larp that they already exist, and then explore what that would mean for a group of people, part of the society, or for our own behaviors and emotions. How this can be done can be seen for instance by the True Color exemplar. Further, larps can be suitable when working with value-based design research, such as value sensitive design, socially responsible design, reflective design, and design for social sustainability, as it is a good way of shifting perspectives, sparking discussions and getting emotional understanding for different concepts and situations. Larps could also be utilized in Research through Design, where one might normally use design ideation methods such as Wizard of Oz (Kelley, 2018), rapid prototyping, use case theatre (Schleicher et al., 2010; Oulasvirta et al., 2003), or object theater - stakeholder drama (Wilde et al., 2017). Larp could be used instead for a deeper first person perspective to try out designs by enacting real world user situations, by roleplaying user reactions and interactions. A key difference between these aforementioned design methods and larp is that larp places a stronger focus toward the emotional, social, and emergent experiences of the participants, rather than having a set outcome or objective. An example of a research project engaging with research through larp could be to create and utilize a sci-fi larp set in a technology-dominated smart city, to gain HCI relevant understanding of how it could be like to live in such a city, using larp as a tool for futuring and speculative design.
@@ 7.1. Benefits of Utilising Larp in HCI
  • Playification - Larp offers many of the benefits of gamification but without strong extrinsic motivations such as badges, scores and leader-boards. Instead, larp motivates through playful exploration and co-creative storytelling.
@@ 7.2. Challenges with Utilising Larp in HCI Research
Despite all these opportunities, there also comes challenges with utilising larp in HCI research. These challenges need to be thought through and reflected upon when making informed decisions as to if and how larp should be utilised. Some of those challenges are more general when designing and facilitating larps, and there is already an abundance of literature on this (Koljonen et al., 2019; Back, 2014; Duus Henriksen et al., 2011), but there are also specific challenges for utilising larp in HCI relevant research, and this is the focus of this section.
@@ 7.3. Handling the Challenges
There is a need to develop robust methods and research practices that are adapted to the design logic of larps. In this, it is vital to learn from fields such as education research, ethnography and anthropology, on how to stage and study situated experiences.
@@
We propose to facilitate more discussions within the HCI community to find new methods for documentation, and develop formats for rigorous mediation of research in relation to larp. There are several practical examples available in the literature that can inform how to structure methods and formats of larps (Waern et al., 2020; Fey et al., 2022), but future research focusing specifically on developing such methods could be helpful for the research community. This discussion needs to also include how design research focused on larp can and should be presented, as the paper format can be ill-suited for situated and embodied experiences such as larp.
@@ 8. Conclusion
  • Larp is for several reasons relevant for the HCI community
@@
Appendix A Table of Larps
Table 5. Table of larps mentioned in the paper. The information is gathered from homepages, and in some cases personal experiences from having attended the larps
Number Name of larp Genre/ Type Short description Year Organizers Country External Link
1 Alkemistens verkstad Online larp/ urban fantasy
An online larp where participants play students at a magical school,
learning alchemistry, connected by magical screens.
2019 LajvVersktaden Sweden https://lajvverkstaden.se/
2 Ancient hours VR larp
VR-larp. The story of Two demi-gods, Love and Memory, held captive
by humans, tortured and pressured into giving up some of their powers.
Conjuring their joint abilities, however, they manage to meet in their
common memories.
2019 Joffe Rydberg Sweden https://alexandria.dk/en/data?scenarie=14058
3 Android Beginning Sci-fi
Bladerunner inspired, philosophical on humanity and androids.
Specially built indoor venue. Nordic larp style, 360°
2022 (campaign) Atropos. Julia Woods. Sweden https://beginnings.atropos.se/
4 Aurum LARP Steampunk/ fantasy
Neo-Victorian Steampunk larp, with dying elf magic, new frontiers
and power-hungry syndicates.
2023 Aurum larp USA https://scifixfantasy.com/events-all/aurum-larp/
5 Belvedere History/ literature larp
Set in an alternative 17th century France, inspired by Three musketeers
and Da Vinci. Technology, prophecies, religion and politics.
2023 RSV Arcana Netherlands https://www.arcana.nl/evenementen/belvedere/
6 Beyond the Neural Horizon Sci-fi
Hackers, where the internet is physically manifested in the rooms.
Specially built indoor larp venue. Nordic larp style, meta techniques.
2019 Atropos Sweden https://alexandria.dk/en/data?scenarie=9900
7 Blodsband Reloded Wasteland
Mad max inspired larp, set in a crazy wasteland future.
Venue: a big sandpit, old industry buildings, and city built of cargo.
2018 (and reruns) Blodsband Reloded Sweden https://lajvhistoria.se/lajv/Blodsband_Reloaded_23.1502
8 Fight like a girl Cyberpunk Dystopian feminist gender war larp 2018 Frida Gamero, Susanne Vejdemo Sweden https://flaglarp.wordpress.com/vision/
9 Fortune & Felicity: A Jane Austen Larp History/ literature larp
Jane Austen inspired, on the drama between passionate love,
necessity of money and the silliness of people. 360° Nordic larp style.
2017 Anna Westerling, Anders Hultman. Sweden https://anna905.wixsite.com/austenlarp
10 Greylight 2142 Cyberpunk
A Cyberpunk world of greed, oppression, and dystopian future.
Political. Indoor venue.
2023 Greylight Germany https://greylight.de/
11 Havenhollow Medieval fantasy Medieval Fantasy, battles and rules. Campaign larp. 2015 (reruns) Underworld Japan https://www.facebook.com/groups/1520452058215307/
12 Insomnia larp Horror Horror larp of investigation of the unknown. 2016 (and reruns) Insomnia LRP UK https://www.facebook.com/insomnialrp/
13 Lovestories by ABBA Musical larp
The players play a band during their last summer tour in the
late 1970s. Focus on love, drama, music and parties.
Emotions are displayed through songs. Chamber/ blackbox larp.
2022 Anna Westerling Sweden https://scenariofestival.se/archive/scenarios-2022/lovestories-by-abba/s-2022/lovestories-by-abba/
14 Mutanternas tidsålder Sci-fi
Childrens summercamp larp about mutants and humans.
Daily larps over a period of 6 weeks
2016 (and reruns) LajvVerkstaden Sweden https://lajvverkstaden.se/
15 Mythodea Fantasy
Huge outdoor fantasy battle larp, around 10 000 participants per larp.
American larp style.
2004 (yearly) Mythodea Germany https://mythodea.de/en/home-en/
16 Nittonhundra Steampunk A steampunk science conference at the turn of the century 2015 Sara Örn Ternstrand Sweden https://steampunkisverige.wordpress.com/category/lajv/
17 Omgiven av idioter Sci-fi/ Convention larp
Homoristic/dystopic larp on social media addiction.
Chamber/ blackbox larp.
2019 Jon Back. Karin Johansson. Sweden https://lajvkonvent.files.wordpress.com/2020/10/broschyr.pdf
18 Terra Incognita Lovecraftian horror/ 1930s
Lovecraftian horror set in the 1930s.
Played in a village and forest. Nordic larp style, 360°
2013
Berättelsefrämjandet
Olle Nyman
Sebastian Utbult
Sweden https://lajvhistoria.se/lajv/Terra_Incognita.45
19 The Austen Experience, A Winter’s Ball
History/Litterature larp.
Cyberpunk, Fusion
Inspired by the works of Jane Austen as well as the tv-show Westworld, and takes place at a Company-owned Austen theme park with AI units and Androids. Location: a castle. Nordic larp style, 360° Atropos Sweden https://austen.atropos.se/
20 The Epic Western Larp, Western Big wild west larp, outdoors and village. 360° 2025 Ultimate Western Spain http://www.ultimewestern.com/en/accueil-en/
21 Turings fråga Future
Philosophical larp set in a non-defined future, were participants try
to pass a test to enter a prestigious academy, by identifying the android
in the group, using different criteria of what it means to be human.
The larp was run in the Arch-Cathedral of Sweden. Nordic larp style.
2013
Frida Gamero
Sara Engström
Johan Dahlberg,
in collaboration with the
Swedish Church.
Sweden https://lajvhistoria.se/lajv/Turings_Fraga.391
22 Voidship Concordia Sci-fi Warhammer 40k -inspired chamber-larp with gamistic elements 2015 Karl Bergström Sweden https://spelkult.se/kalender/event/voidship-concordia/
23 Zero Sci-fi
Players play a crew aboard a spaceship, leaving the dying earth behind.
A technology enhanced spaceship venue. Futuring, political,
Nordic larp style, 360°
2023 Not Only Larp Spain https://zero.notonlylarp.com/
24 Last Will Dystopian Future
A larp aimed at looking at modern slavery through the lens of a fictional future.
A fading human dignity in a world run by money and consumption in which
people can be bought and sold as commodities.
Nordic larp style, 360°, combined with meta techniques
2014,2015
Ursula
Annica Strand
Frida Gamero
Sofia Stenler
Sweden https://www.foreningenursula.se/last-will/
25 Odysseus Sci-fi
Survivors from a destroyed space colony try to find the home of their ancestors.
2019, 2024 Illusiary Finland https://www.odysseuslarp.com/
26 Mission Together Sci-fi
A space mission full of power struggles and cultural differences
between people from different planets.
2020 Not Only Larp Spain https://notonlylarp.com/